OpenStudy (anonymous):

Find the Domain and Range and sketch the graph. 1. f(x)=√(x-1)-1 2. f(x)= 3 - √2-x 3. f(x)=√4 -(x+2)^2 4. f(x)=3 + √x^2+1 5. f(x)=√x-1 -1

OpenStudy (anonymous):

Domain is the set of x values for the equation and the range is the set of y values for the equation. Example 1) f(x) = sqrt (x - 1) - 1 Domain: x >= 1 because if x is smaller than 1, you would get the sqrt of a negative number and that's not the graph you are going to sketch. Range: f >= -1. Substitute the endpoints of the domain into the equation and most likely get the answer. f(1) = sqrt (1 - 1) - 1 = -1.